# Efficient detection of chaos through the computation of the Generalized Alignment Index (GALI) by the multi-particle method

### v1

@@ Acknowledgments
The authors are grateful to Henok T. Moges for useful discussions. We thank the Centre for High Performance Computing (CHPC) of South Africa for providing the computational resources. M.H. acknowledges support from the National Research Foundation (NRF) of South Africa (Grant No. 129630).
Appendix A Pseudo code for the computation of the GALI using the MPM
@@
    Equations of motion for the system
@@
    Initial conditions to integrate
@@
    The order $k$ of GALI to compute
@@
    Initial orthogonal separation vectors, $w_{i}(0)$
@@
    Separation vector size $d_{0}$
@@
    Time step $\delta t$
@@
    Final time $t_{f}$
  Initialize nearby particles $xn_{i}$ from deviations
@@
Set the next time we renormalize separation vectors to $\tau$
@@
Iterate forward from time $t=0$ and compute GALI
@@
In the above pseudo-code, the SVD denotes the single value decomposition routine, which has demonstrated its efficiency for the computation of GALIs.
Appendix B Initial conditions for $\beta-$FPUT lattice with $N=5$
The initial conditions for the orbits used to compute the MLE and the GALI in Figs. 10, 11 and 12 are as follows. The regular orbit is given by
$\begin{split}x_{1}&=1.03003\\ x_{2}&=0\\ x_{3}&=-1.04003\\ x_{4}&=0\\ x_{5}&=1.04003\\ p_{1}&=0.29284\\ p_{2}&=p_{3}=p_{4}=p_{5}=0,\end{split}$ (42)
while the chaotic one reads
$\begin{split}x_{1}&=-1.151344372237934177\\ x_{2}&=-0.000000000000007443\\ x_{3}&=1.151344372237925295\\ x_{4}&=0.000000000000006056\\ x_{5}&=-1.151344372237918634\\ p_{1}&=1.502757415151290132\\ p_{2}&=0.000000000000014786\\ p_{3}&=-1.502757415151279696\\ p_{4}&=-0.000000000000017280\\ p_{5}&=1.502757415151272147.\end{split}$ (43)
Appendix C Reliable regions of GALI computed using the MPM for various $\beta-\mbox{FPUT}$ lattice sizes
In Section 4.3 we presented results for the computation of GALI using an FPUT lattice of five particles. Here for completeness we show the reliable region of the same GALIs calculated for chaotic initial conditions, of the order $2,4,6,$ and 8, computed using more degrees of freedom. Figure 13 shows the final value of the GALIs computed with the MPM (orange circles) and the VM (blue squares) using $N=11,N=101$, and $N=501$ particles. Here we see the same pattern as with $N=5$, where the optimal deviation vector magnitude is around $d_{0}\approx 10^{-8}$ in all cases, and an approximate if asymmetric V-shape is visible in the values of GALI computed through the MPM. As the order of the GALI increases the asymmetry of this shape increases, likely due to error terms arising from longer-time effects than predicted by the short-time theory. These results confirm that the overall findings hold true even for relatively many degrees of freedom.
Figure 13: $\bm{\beta-}$FPUT chain. Diagram of reliable regions for $\mbox{GALI}_{k}$ in the bFPUT computed by the multi-particle method for various lattice sizes. The total norm of the deviation vector is varied between $10^{-14}$ and $10^{-2}$. Columns correspond to values of $N=11$, $101$, and $501$ respectively, while rows correspond to GALI 2, 4, 6 and 8. The blue squares are the values of $\mbox{GALI}_{k}$ computed through the variational method, while the orange circles are the data computed via the multi-particle method. The GALI value was measured at time $t=700$ ($N=11$), $t=700$ ($N=1500$), and $t=7000$ ($N=501$). The energies are chosen just above the critical energies $\mathcal{H}_{11}^{c}=1.98$, $\mathcal{H}^{c}_{101}=1.515$ and $\mathcal{H}_{501}^{c}=1.502$ respectively for $N=11$, $101$ and $501$.

### v2

@@ Acknowledgments
We are grateful to Henok Moges for useful discussions. We also thank the Centre for High Performance Computing (CHPC) of South Africa for providing computational resources for this project. M.H. acknowledges support from the National Research Foundation (NRF) of South Africa (Grant No. 129630). The authors would like to express their gratitude to the anonymous reviewers for their comments and suggestions, which helped improve the clarity of this manuscript.
Appendix A Pseudo-code for the computation of the generalized alignment index using the multi-particle method
@@
    Equations of motion of the system.
@@
    Initial conditions of the tested orbit.
@@
    The order $k$ of the computed GALI.
@@
    Initial orthogonal separation vectors, $w_{i}(0)$, $i=1,2,\ldots,k$.
@@
    Separation vector size $d_{0}$.
@@
    Time step $\delta t$.
@@
    Final integration time $t_{f}$.
  Create initial conditions $xn_{i}$ for the nearby orbits from separation vectors.
@@
Set the first renormalization time.
@@
Forward integration from time $t=0$ and computation of the GALIk.
@@
In the above pseudo-code, the SVD denotes the ‘single value decomposition’ method, which has been used in [60] to efficiently compute GALIs.
Appendix B Initial conditions of orbits of the $\beta$-Fermi-Pasta-Ulam-Tsingou lattice with $N=5$
The initial conditions of the orbits of the $\beta$-FPUT model (39) with $N=5$ used to compute the MLE and the GALI in Figs. 10, 11 and 12 are
$\begin{split}x_{1}&=1.03003,\\ x_{2}&=0,\\ x_{3}&=-1.04003,\\ x_{4}&=0,\\ x_{5}&=1.04003,\\ p_{1}&=0.29284,\\ p_{2}&=p_{3}=p_{4}=p_{5}=0,\end{split}$ (42)
for the regular orbit, and
$\begin{split}x_{1}&=-1.151344372237934177,\\ x_{2}&=-0.000000000000007443,\\ x_{3}&=1.151344372237925295,\\ x_{4}&=0.000000000000006056,\\ x_{5}&=-1.151344372237918634,\\ p_{1}&=1.502757415151290132,\\ p_{2}&=0.000000000000014786,\\ p_{3}&=-1.502757415151279696,\\ p_{4}&=-0.000000000000017280,\\ p_{5}&=1.502757415151272147,\end{split}$ (43)
for the chaotic one.
Appendix C Reliable regions of the generalized alignment index computed using the multi-particle method for various $\beta$-Fermi-Pasta-Ulam-Tsingou lattice sizes
In Sect. 4.3 we presented results for the computation of the GALIs using a $\beta$-FPUT lattice of $N=5$ oscillators. Here, for completeness sake, we discuss the reliable region of some GALI computations for chaotic orbits of the $\beta$-FPUT model with more degrees of freedom. The considered chaotic orbits are located in the neighborhood of the relevant unstable SPO1 trajectory for energies just above the first energy destabilization of the SPO1 happening at $\mathcal{H}_{11}^{c}=1.98$, $\mathcal{H}^{c}_{101}=1.515$ and $\mathcal{H}_{501}^{c}=1.502$ respectively for $N=11$, $101$ and $501$. In Fig. 13 we present the final value of the GALI2 (first, top row), GALI4 (second row), GALI6 (third row) and GALI8 (fourth, bottom row) computed by the MPM (red circles) and the VM (blue triangles) for $N=11$ (left column), $N=101$ (middle column), and $N=501$ (right column) oscillators. Here we see again the same pattern which was observed for the $N=5$ [Fig. 12], with the optimal deviation vector magnitude being $d_{0}\approx 10^{-8}$ for all cases, along with the formation of a $V-$shaped feature, which becomes more asymmetric as the order of the GALI increases. These results confirm that the main findings of our theoretical analysis for the determination of the optimal $d_{0}$ value, are also valid for Hamiltonian systems with many degrees of freedom.
Figure 13: $\bm{\beta}$-FPUT chain. Dependence of the GALI2 (top row), GALI4 (second row), GALI6 (third row) and GALI8 (bottom row) values on the norm $d_{0}$ of the initial deviation vectors for chaotic orbits of the $\beta$-FPUT lattice model (39) $N=11$ (left column), $N=101$ (middle column), and $N=501$ (right column) oscillators, when the the MPM (red circles) and the VM (blue triangles) are used. In all cases the norm $d_{0}$ is varied between $10^{-14}$ and $10^{-2}$. The GALI values are computed at time $t=700$ ($N=11$), $t=700$ ($N=101$), and $t=7000$ ($N=501$). The energies of the considered orbits are chosen just above the critical energies $\mathcal{H}_{11}^{c}=1.98$, $\mathcal{H}^{c}_{101}=1.515$ and $\mathcal{H}_{501}^{c}=1.502$ respectively for $N=11$, $101$ and $501$.